# Scaling Offline Model-Based RL via Jointly-Optimized World-Action Model Pretraining

### v1

@@ 5.2 How does JOWA perform on multi-games in low-data regime?
Game Random Human MTBC MGDT EDT SQL JOWA
120M 200M 200M 80M 40M 70M 150M
Assault 222.4 742.0 1203.5 1741.5 1915.5 2292 1578.3 1733.9 2302
Atlantis 12850 29028.1 37812.5 2565750 2108166.7 49100 41662.5 570862.5 2690387.5
BeamRider 363.9 16926.5 786 6011.3 4149.8 7023.5 880.3 2547.4 3498
Berzerk 123.7 2630.4 655 444.5 350 312.5 395 441.9 739
Carnival 0 3800 5560 2610 3678.8 1940 4685 4070 5316
Centipede 2090.9 12017 4046.6 4604 3389.9 3650.3 5669.3 4475.6 4677
ChopperCommand 811.0 7387.8 256.3 3300.8 3443.8 853.8 3118.8 2568.8 3812.5
DemonAttack 152.1 1971 2611.3 6549.4 3455.7 7936.5 1233.8 4584.4 3547.8
NameThisGame 2292.3 8049 6045 6610.5 7060 6461.7 7335.6 12706.9 11421
Phoenix 761.4 7242.6 1446.9 5120.5 5320 4961.7 1608.8 5065 5348
Seaquest 68.4 42054.7 120 2720 3160.4 650 1033.1 1490 2725
SpaceInvaders 148 1668.7 605 742.5 513.8 714.4 694.7 969.1 744.7
StarGunner 664.0 10250 1493.8 8625.5 9550 4728.6 5737.5 21231.3 18150
TimePilot 3568 5229.2 762.5 3866.7 2812.5 4072.7 3268.8 3831.3 3669
Zaxxon 32.5 9173.3 0 462.5 325.5 0 0 225 2163
#Superhuman 0 N/A 4 3 4 3 3 4 6
Median HNS 0.000 1.000 0.197 0.391 0.400 0.387 0.360 0.390 0.456
IQM HNS 0.000 1.000 0.329 0.498 0.502 0.420 0.371 0.476 0.789
We summarize our main results in Table 2. This table shows the performance of JOWA alongside all best performing sizes of baselines trained with 10% subsampled dataset. MTBC, MGDT, and EDT represent (conditional) behavior cloning methods, while Scaled-QL and JOWA represent Q-learning methods. Despite the constraints on the amount of data, JOWA achieves superior performance with comparable or fewer parameters than baselines. Specifically, JOWA-150M attains an IQM human-normalized score of 78.9% in low-data regime, surpassing MGDT-200M (49.8%) and EDT-200M (50.2%), despite these two models having more parameters and utilizing data augmentation during pretraining. JOWA-70M achieves an IQM human-normalized score of 47.6%, outperforming Scaled-QL-80M (42.0%). JOWA maintains its performance advantage when comparing median human-normalized scores. These results demonstrate JOWA’s sample efficiency in learning from heterogeneous offline data.
@@ 5.3 How does JOWA scales with model size?
Figure 2: Scaling trends for different algorithms on training set games.
@@
Game MTBC MGDT EDT SQL JOWA JOWA-150M
120M 200M 200M 80M 150M (scratch)
Mean 0.164 0.422 0.430 0.360 0.647 0.196
Median 0.215 0.354 0.325 0.284 0.615 0.173
IQM 0.205 0.377 0.380 0.355 0.647 0.181
@@ 5.4 Can JOWA sample-efficiently transfer to new games?
Rapid adaptation to downstream tasks is a natural and well-motivated benefit of pretraining. In this section, we study how the pretrained world-action model enables rapid and sample-efficient fine-tuning on new games. To investigate this question, we fine-tune pretrained agents on 5 held-out games using uniformly subsampled 5k expert-level transitions (last 20% of DQN-Replay (Agarwal et al., 2020)) per game as the benchmark. These tiny amounts of transitions corresponding to approximately 4 trajectories from DQN-Replay per fine-tuned game on average, which is similar to the settings of few-shot learning and extremely challenging. For comparison, we fine-tune the largest agents of baselines alongside a no-pretrain baseline, JOWA-150M (scratch) trained on each held-out game from scratch. All pretrained agents are fine-tuned for 10k steps while JOWA-150M (scratch) is trained for 100k gradient steps. Detailed fine-tuning protocol is shown in Appendix E.2. We report results in terms of DQN-normalized score, following Lee et al. (2022); Kumar et al. (2023).
We summarize the fine-tuning results in Table 3. This table shows that the fine-tuned JOWA-150M attains 64.7% IQM DQN-normalized score across 5 held-out games, outperforming baselines by 34.7% on averange. These results demonstrates that the learned general-purpose representation and decision-making of JOWA helps rapid and sample-efficient transfer to novel games.
@@ 5.5 Ablation Study
In this section, we conduct a series of controlled ablation studies to evaluate the impact of key design choices in JOWA, including planning, different training losses, and architecture designs. These experiments aim to provide empirical evidence to support our design decisions and offer valuable insights for future research in this domain.
Table 4: Performance on the 15 Atari games for 3 variants of JOWA evaluated with or without planning reported in terms of the IQM human-normalized score.
@@
We evaluate our 3 variants models with or without planning on the pretrained games and report the IQM aggregated human-normalized score across 15 games in Table 4. Observe that the addition of planning improves 13.5% performance on average and 28.1% on maximum for JOWA-150M. Even without planning, JOWA still demonstrates scaling trends with the growth of parameter size, but at a slower scaling rate.
@@
Table 5: The mean, median, and IQM human-normalized score on the 6 Atari games for various training choices. See Tabel 18 in Appendix F for raw scores.
Mean HNS 1.183 0.613 0.917 0.293 0.189 1.209 1.026 0.964 0.448
Median HNS 1.078 0.696 0.489 0.304 0.118 0.975 0.921 0.721 0.452
IQM HNS 1.123 0.637 0.659 0.307 0.126 1.049 0.931 0.824 0.464
@@
Observe that the $\operatorname{sg}(\mathcal{L}_{\text{action}})$ and MSE loss fail to train qualified multi-game agents. We empirically observe that agents trained with MSE TD-loss always over-optimize the CQL regularization term regardless of the value of coefficient $\alpha$ (tested with $\alpha\in\{0.01,0.05,0.1\}$), resulting in extreme Q-values for in-domain state-actions and low Q-values for OOD actions. The fail of $\operatorname{sg}(\mathcal{L}_{\text{action}})$ underscores the importance of joint-optimization in the world-action model. Comparing the original loss with the no $\mathcal{L}_{\text{world}}$ configuration demonstrates that scaling model-based multi-game RL is more efficient than model-free RL. Overall, the original loss outperforms all variant losses by a large margin, indicating the effectiveness of every loss terms and their combination proves essential for the overall success of JOWA.
Ensemble of Q-value heads.
We run JOWA-40M with ensembled Q-heads, where the Q-distribution is a weighted sum of distributions from multiple Q-heads. The original configuration uses a single Q-head. In this experiment, we set the number of multi Q-heads to 3 and compare 2 ensemble approaches: equal weights or random weights like REM (Agarwal et al., 2020). We report the results in the forth main column of Table 5. Although the equal weighted ensemble slightly outperforms random weights, we have not observed significant improvements of multi Q-heads over single Q-head. We leave the better way of distributional-Q ensemble for future work.
Effects of task embedding.
@@
By default, we do not use planning or synthetic data during pretraining due to the significant increase in training time (approximately 5 times slower, extending training to over a month). For this ablation study, we enable synthetic data in pretraining for JOWA-40M and train on 6 games for 1M steps with the same batch size and all other hyperparameters, which takes 31 days. Specifically, we sample batch of 4-frames segments and interact in the imagined MDP with $\epsilon$-greedy policy to synthesis the last 4 steps. Then we optimize the world-action model with half batch of real data and half batch of synthetic data using COMBO (Yu et al., 2021) loss in Equation (4) as $\mathcal{L}_{\text{action}}$. The results are shown in the last main column of Table 5.
Suprisingly, we observe negative gains from incorporating synthetic data. We hypothesize twofold reasons: (i) accumulation of inference errors over steps causes synthetic steps to deviate significantly from the ground-truth distribution, (ii) COMBO’s over-penalization of Q-values for unseen state-action pairs results in overly conservative agents. Due to the unbearable training time, further investigation into more effective methods for utilizing synthetic data in multi-game RL pretraining is left for future work.
@@ 6 Conclusion
In this work, we introduce JOWA: Jointly-Optimized World-Action model, a single offline model-based RL agent capable of playing multiple Atari games. JOWA uses a shared transformer backbone for both the world modeling and the Q-value criticism, enabling joint optimization. We propose a parallelizable planning algorithm to consistently identify the optimal policy during inference. As we hoped, by increasing model parameters, JOWA unlocks scaling trends in performance and exceed prior large-scale offline RL methods in multi-games regime. Furthermore, by training a large-capacity model on a diverse set of games, we show that JOWA can sample-efficiently adapt to novel games, leveraging its generalizable world model for planning. Our ablation studies validate the efficacy of joint optimization and our planning method, while also demonstrating that scaling offline model-based RL is more efficient than offline model-free RL. To facilitate future research, we will release all training and evaluation codes, along with pretrained model checkpoints.
@@
Due to the computation resources required, we could not experiment on full Atari 2600 games with complete datasets, as this would take approximately two months of training time. Additionally, while using synthetic data to train agents is common in single-task online model-based RL algorithms, we observe it yields negative gains in multi-game RL pretraining. Therefore, how to effectively use synthetic data for multi-task pretraining is an interesting direction for future work.
@@ Appendix A Proof of Equation (11)
For Equation (16), we derive the optimal policy $\pi$ is the greedy policy selecting actions with the greatest Q*-value:
@@ Appendix D Games
We select 20 Atari games maintaining the difficulty distribution of full Atari 2600 games defined by Gulcehre et al. (2020), which includes 9 easy games, 9 medium games, and 2 hard games. We use 15 out of 18 games for training and the remaining 5 for OOD generalization experiments. The 15 training games are: Phoenix, Centipede, SpaceInvaders, Carnival, NameThisGame, Assault, Atlantis, DemonAttack, BeamRider, ChopperCommand, Seaquest, TimePilot, StarGunner, Berzerk, Zaxxon. The 5 held-out games are: Pong, Robotank, YarsRevenge, Gravitar, MsPacman. Details about the size of action spaces and game difficulties are shown in Table 6.
@@ E.1.1 JOWA
where $E,D,\mathcal{E}$ are encoder, decoder, and embedding table respectively. $\mathrm{sg}(\cdot)$ is the stop-gradient operator. The last term is the perceptual loss (Johnson et al., 2016). We list the hyperparameters of VQ-VAE in Table 7 and 8. After the first stage of training, the VQ-VAE is frozen.
@@
In addition to the vocabulary embedding and position embedding, we add a learnable task embedding for observation tokens and action tokens respectively. Our transformer are based on minGPT [^2] https://github.com/karpathy/minGPT with FlashAttention (Dao et al., 2022) for acceleration. The hyperparameters of JOWA’s transformer backbone are listed in Table 9 and 10.
@@
Model Layers Hidden size Heads Params
JOWA-40M 4 512 8 37.87M
JOWA-70M 6 768 12 72.11M
JOWA-150M 12 768 12 146.54M
@@ E.1.5 Scaled-QL
For fair comparison, all methods are trained with the same batch size of 512 for 1.75M gradient steps.
@@ E.2 Fine-tuning protocol
We uniformly draw 5k transitions from expert level DQN-Replay (Agarwal et al., 2020) (last 20% of the original dataset) for each held-out game. Each game was fine-tuned separately to measure the model’s transfer performance for a fixed game. we fine-tuned all methods using a batch size of 32 and learning rate of 0.00005 for 10k gradient steps. For reporting results, we report the performance of the agent snapshot that obtain the highest score during fine-tuning.
For JOWA in the second fine-tuning stage, we set both the planning horizon and the beam width to 2 for all fine-tuning experiments. Thus we sample batch of 6-steps segments, using planning algorithm to synthesis the last 2 steps. Then we update JOWA with 3/4 batch (24) of real data and 1/4 batch (8) of synthetic data using COMBO loss rather than CQL loss. For other baselines, we enable random cropping and random rotation for data augmentation.

### v3

@@ 5.3 How does JOWA perform on multi-games in low-data regime?
Game Random Human MTBC MGDT EDT SQL FICC JOWA
120M 200M 200M 80M 85M 40M 70M 150M
Assault 222.4 742.0 1203.5 1741.5 1915.5 2292 925.9 1578.3 1733.9 2302
Atlantis 12850 29028.1 37812.5 2565750 2108166.7 49100 86250 41662.5 570862.5 2690387.5
BeamRider 363.9 16926.5 786 6011.3 4149.8 7023.5 6822 880.3 2547.4 3498
Berzerk 123.7 2630.4 655 444.5 350 312.5 400 395 441.9 739
Carnival 0 3800 5560 2610 3678.8 1940 2820 4685 4070 5316
Centipede 2090.9 12017 4046.6 4604 3389.9 3650.3 3742.2 5669.3 4475.6 4677
ChopperCommand 811.0 7387.8 256.3 3300.8 3443.8 853.8 2835.6 3118.8 2568.8 3812.5
DemonAttack 152.1 1971 2611.3 6549.4 3455.7 7936.5 5806.4 1233.8 4584.4 3547.8
NameThisGame 2292.3 8049 6045 6610.5 7060 6461.7 6236 7335.6 12706.9 11421
Phoenix 761.4 7242.6 1446.9 5120.5 5320 4961.7 3814.5 1608.8 5065 5348
Seaquest 68.4 42054.7 120 2720 3160.4 650 1760 1033.1 1490 2725
SpaceInvaders 148 1668.7 605 742.5 513.8 714.4 641.2 694.7 969.1 744.7
StarGunner 664.0 10250 1493.8 8625.5 9550 4728.6 4936.4 5737.5 21231.3 18150
TimePilot 3568 5229.2 762.5 3866.7 2812.5 4072.7 4166.7 3268.8 3831.3 3669
Zaxxon 32.5 9173.3 0 462.5 325.5 0 312.5 0 225 2163
#Superhuman 0 N/A 4 3 4 3 3 3 4 6
Median HNS 0.000 1.000 0.197 0.391 0.400 0.387 0.390 0.360 0.390 0.456
IQM HNS 0.000 1.000 0.329 0.498 0.502 0.420 0.433 0.371 0.476 0.789
We summarize our main results in Table 2. This table shows the performance of JOWA alongside all best performing sizes of baselines trained with 10% subsampled dataset. MTBC, MGDT, and EDT represent (conditional) behavior cloning methods, while Scaled-QL, FICC, and JOWA represent Q-learning methods. Despite the constraints on the amount of data, JOWA achieves superior performance with comparable or fewer parameters than baselines. Specifically, JOWA-150M attains an IQM human-normalized score of 78.9% in low-data regime, surpassing MGDT-200M (49.8%) and EDT-200M (50.2%), despite these two models having more parameters and utilizing data augmentation during pretraining. JOWA-70M achieves an IQM human-normalized score of 47.6%, outperforming Scaled-QL-80M (42.0%) and FICC-85M (43.3%). JOWA maintains its performance advantage when comparing median human-normalized scores. These results demonstrate JOWA’s sample efficiency in learning from heterogeneous offline data.
Next, we present experiments highlighting JOWA’s scaling and generalization capabilities. These two key properties derived from multi-game pretraining align with recent advancements in SL.
@@ 5.4 How does JOWA scales with model size?
Figure 2: Scaling trends for different algorithms on the training set games.
@@
DNS MTBC MGDT EDT SQL FICC JOWA JOWA-150M
120M 200M 200M 80M 85M 40M 70M 150M (scratch)
Mean 0.164 0.422 0.430 0.360 0.543 0.504 0.576 0.647 0.196
Median 0.215 0.354 0.325 0.284 0.565 0.512 0.715 0.615 0.173
IQM 0.205 0.377 0.380 0.355 0.575 0.498 0.603 0.647 0.181
@@ 5.5 Can JOWA sample-efficiently transfer to new games?
Rapid adaptation to downstream tasks is a natural and well-motivated benefit of pretraining. To investigate this question, we fine-tune pretrained agents on 5 held-out games using uniformly subsampled 5k expert-level transitions (from last 20% of DQN-Replay (Agarwal et al., 2020)) per game as the benchmark. These tiny amounts of transitions corresponding to approximately 4 trajectories from expert-level DQN-Replay per fine-tuned game on average, which is similar to the settings of few-shot learning and is extremely challenging. For comparison, we fine-tune the largest agents of baselines alongside a no-pretrained baseline, JOWA-150M (scratch) trained on each held-out game from scratch. All pretrained agents and JOWA-150M (scratch) are fine-tuned for 50k and 500k gradient steps, respectively. Detailed fine-tuning protocol is shown in Appendix E.2. We report results in terms of DQN-normalized score (DNS), following Lee et al. (2022), in Table 3.
The results show that the fine-tuned JOWA-150M attains 64.7% IQM DNS across 5 held-out games, outperforming baselines by 69.9% on average. Moreover, comparing the performance of 3 JOWA variants, the scaling trend on IQM DNS still holds for novel games after fine-tuning. These results underscore JOWA’s capacity for rapid and sample-efficient transfer to novel games, highlighting the efficacy of its learned general-purpose representation and decision-making capabilities.
@@ 5.6 Ablation Study
In this section, we conduct a series of ablation studies to evaluate the impact of key design choices in JOWA, including planning, different training losses, and usage of synthetic data. More experiments can be found in Appendix F. These experiments aim to provide empirical evidence to support our designs and offer valuable insights for future research in this domain.
Table 4: Performance on the 15 Atari games for 3 variants of JOWA evaluated with or without planning reported in terms of the IQM HNS.
@@
Table 5: Results on 7 games for JOWA-150M evaluated with different planning algorithms. See Table 19 for raw scores.
w/o planning MCTS Ours
FPS 10.8 0.12 1.26
Mean HNS -2% 22.1% 37.8%
IQM HNS 3% 13.4% 23.7%
@@
We evaluate our 3 variants models with or without planning and report the IQM HNS across 15 pretrained games in Table 4. Observe that the addition of planning improves 13.5% performance on average and 28.1% on maximum for JOWA-150M. Even without planning, JOWA still exhibits scaling trends with increasing model size, but at a slower rate.
To further demonstrate the efficiency of our planning algorithm, we compare it with MCTS. Because Muzero-style (Schrittwieser et al., 2020) MCTS requires access to the $V$-function and policy networks $\pi$ while JOWA only estimates the optimal $Q$-function, we use $\max_{a}Q(s,a)$ as the $V$-value $V(s)$ and use an energy-based policy to compute the action probability, i.e., $\pi(\cdot|s)=\operatorname{softmax}(Q(s,\cdot)/t)$, where $t$ is the temperature. We conduct a grid search on the implementation choices and hyperparameters for MCTS and show the details in Appendix F. We report the frames per second (FPS) and HNS on 7 games, where bare JOWA-150M performs poorly, in Table 5. Observe that ours not only is 10$\times$ faster than MCTS but also exceeds MCTS by 71.0% on Mean HNS. Moreover, we empirically observe that MCTS is highly sensitive to the temperature and max depth of the tree while the long execution time makes it inconvenient to tune hyperparameters.
@@
Table 6: The mean, median, and IQM human-normalized score on the 6 Atari games for various training choices. See Tabel 20 in Appendix F for raw scores.
Game Origin Different training losses Synthetic data
No CQL No $\mathcal{L}_{\text{world}}$ $\operatorname{sg}(\mathcal{L}_{\text{action}})$ MSE in pretraining
Mean HNS 1.183 0.613 0.917 0.293 0.189 0.448
Median HNS 1.078 0.696 0.489 0.304 0.118 0.452
IQM HNS 1.123 0.637 0.659 0.307 0.126 0.464
@@
Observe that the $\operatorname{sg}(\mathcal{L}_{\text{action}})$ and MSE loss fail to train qualified multi-game agents. We empirically observe that agents trained with MSE TD-loss always over-optimize the CQL regularization term regardless of the value of coefficient $\alpha$ (tested with $\alpha\in\{0.01,0.05,0.1\}$), resulting in extremely high Q-values for in-domain state-actions and low Q-values for OOD actions. The fail of $\operatorname{sg}(\mathcal{L}_{\text{action}})$ underscores the importance of joint-optimization in the world-action model. Comparing the original loss with the no $\mathcal{L}_{\text{world}}$ configuration demonstrates that scaling model-based multi-game RL is more efficient than model-free RL. Additionally, the CQL conservatism regularizer is also necessary in multi-game pretraining. Overall, the original loss outperforms all variant losses by a large margin, indicating the effectiveness of every loss terms.
@@
By default, we do not employ planning or synthetic data during pretraining due to the significant increase in training time (approximately 10$\times$ slower, extending training to months). For this ablation study, we enable synthetic data in pretraining for JOWA-40M and train on 6 games for 1M steps, which takes 31 days. Specifically, we sample batch of 4-step segments and interact in the imagined MDP with $\epsilon$-greedy policy to synthesis the last 4 steps. Then we optimize JOWA with half real data and half synthetic data using COMBO loss as $\mathcal{L}_{\text{action}}$. The results are shown in the last main column of Table 6.
Suprisingly, we observe negative gains from the usage of synthetic data. We hypothesize twofold reasons: (i) accumulation of inference errors over steps causes later synthetic steps to deviate significantly from the ground-truth distribution, (ii) COMBO’s over-penalization of Q-values for unseen state-action pairs results in overly conservative agents. Due to the unbearable training time, further investigation into more effective methods of utilizing synthetic data is left for future work.
@@ 6 Conclusion
In this work, we introduce JOWA: Jointly-Optimized World-Action model, a single offline model-based RL agent capable of playing multiple Atari games. JOWA uses a shared transformer backbone for both the world modeling and the Q-value criticism, enabling joint optimization. We propose a provably efficient and parallelizable planning algorithm to consistently identify the optimal policy during inference. As we hoped, by increasing model parameters, JOWA unlocks scaling trends in performance and exceed prior large-scale offline RL methods in multi-games regime. Furthermore, by training a large-capacity model on a diverse set of games, we show that JOWA can sample-efficiently adapt to novel games, leveraging its generalizable world model for planning. Our ablation studies validate the efficacy of joint optimization and our planning method, while also demonstrating that scaling offline model-based RL is more efficient than offline model-free RL. To facilitate future research, we will release all training and evaluation codes, along with pretrained model checkpoints.
@@
Due to the computation resources required, we could not experiment on full Atari 2600 games with complete datasets (Agarwal et al., 2020), as this would take 2 months for training. Additionally, while using synthetic data to train agents is common in single-task online model-based RL, we observe it yields negative gains in multi-game RL pretraining. Therefore, how to effectively use synthetic data for multi-task pretraining is an interesting direction for future work. Lastly, although we observe the scaling trend on IQM aggregated scores, the raw scores in some games are against this trend, i.e.,increasing model parameters leads to decreasing performance in specific games. We hypothesis that this inconsistence is an inherent problem of offline RL. Unfortunately, to our knowledge, neither increasing model size nor extending training iterations can solve this issue.
Acknowledgments
We are grateful to the anonymous reviewers and editors for their insightful suggestions. This work was partially supported by the National Science and Technology Major Project (2022ZD0117102), the National Natural Science Foundation of China under Grants 62271485 and 62303462, Beijing Natural Science Foundation under grant L241016 and L233005, Chongqing Transportation Technology Project (CQJT-CZKJ2024-04), Sichuan Key-Area Research and Development Program 2024YFHZ0011, the Provincial Key Research and Development Program of Zhejiang (project number: 2022C01129), Ningbo International Science and Technology Cooperation Project (2023H020).
@@ Appendix A Proof of Equation (10)
For Equation (15), we derive the optimal policy $\pi$ is the greedy policy selecting actions with the greatest Q*-value as follows:
@@ Appendix D Games
We select 20 Atari games maintaining the difficulty distribution of full Atari 2600 games defined by Gulcehre et al. (2020), which includes 9 easy games, 9 medium games, and 2 hard games. We use 15 out of 20 games for training and the remaining 5 for OOD generalization experiments. The 15 training games are: Phoenix, Centipede, SpaceInvaders, Carnival, NameThisGame, Assault, Atlantis, DemonAttack, BeamRider, ChopperCommand, Seaquest, TimePilot, StarGunner, Berzerk, Zaxxon. The 5 held-out games are: Pong, Robotank, YarsRevenge, Gravitar, MsPacman. Details about the size of action spaces and game difficulties are shown in Table 7.
@@ E.1.1 JOWA
where $E,D,\mathcal{E}$ are encoder, decoder, and embedding table respectively. $\mathrm{sg}(\cdot)$ is the stop-gradient operator. The last term is the perceptual loss (Johnson et al., 2016). We list the hyperparameters of VQ-VAE in Table 9 and 9. After the first stage of pretraining, the VQ-VAE is frozen.
@@
In addition to the vocabulary embedding and position embedding, we add a learnable task embedding for observation tokens and action tokens respectively. Our transformer are based on minGPT [^2] https://github.com/karpathy/minGPT with FlashAttention (Dao et al., 2022) for acceleration. The hyperparameters of JOWA’s transformer backbone are listed in Table 11 and 11.
@@
Model Layers Hidden size Heads
JOWA-40M 4 512 8
JOWA-70M 6 768 12
JOWA-150M 12 768 12
@@ E.1.5 Scaled-QL
For fair comparison, all methods are trained with the same batch size of 512 for 1.75M gradient steps. For reporting results, we report the performance of the agent at the end of pretraining.
@@ E.2 Fine-tuning protocol
We uniformly draw 5k transitions from expert-level DQN-Replay (Agarwal et al., 2020) (last 20% of the original dataset) for each held-out game. Each game was fine-tuned separately to measure the model’s transfer performance for a fixed game. we fine-tuned all methods using a batch size of 32 and learning rate of 0.00005 for 50k gradient steps. For reporting results, we report the performance of the agent snapshot that obtain the highest score during fine-tuning.
For JOWA in the second fine-tuning stage, we set both the planning horizon and the beam width to 2 for all fine-tuning experiments. Thus we sample batch of 6-steps segments, using planning algorithm to synthesis the last 2 steps. For other baselines, we enable random cropping and random rotation for data augmentation.